the equation 7x-8a=3x+24a for x
Answer by checkley77(12844)

the equation 7x-8a=3x+24a for x
7x-3x=24a+8a
4x=32a
x=32a/4
x=8a ans.
Proof:
7*8a-3*8a=24a+8a
56a-24a=32a
32a=32a